OpenStudy (anonymous):

What is the GCF of the following expression: 24z^7-16z^4

OpenStudy (anonymous):

I got 6z^4, but I haven't done a problem like that in a little while. Can anyone back me up or prove me wrong? lol

OpenStudy (raden):

24 = 2^3 * 3 16 = 2^4 The GCF of 24 and 16 = 2^3 = 8 while the GCF of z^4 and z^7 is z^4 therefore, the GCF of them = 8z^4

OpenStudy (anonymous):

Oh right. Added the groups of 2 instead of multiplying. Haha. so 8 is right, not 6.

OpenStudy (raden):

hehe ;P

OpenStudy (anonymous):

thank you man can i throw another one out

OpenStudy (anonymous):

identify one of the two solutions for : x^2-15x-54=0

OpenStudy (anonymous):